3-78.Thickness measurements of a coating process are made to the nearest hundredth of a millimeter. The thickness measurements are uniformly distributed with values 0.15, 0.16. 0.17, 0.18, and 0.19. Determine the mean and variance of the coating thickness for this process
Answer by Boreal(15235)

The mean is the middle of this uniform function that has a=0.15 and b=0.19, so that is (a+b)/2=0.17
the variance is (b-a)^2/12
=0.04^2/12=0.0001333